OpenStudy (anonymous):

If gcd(a,b)=1 and c|a, then gcd(b,c)=1

OpenStudy (helder_edwin):

since g.c.d.(a,b)=1 then there exist integers x0 and y0 such that \[ \large 1=ax_0+by_0 \] also since \(c\mid a\) then there exists k integer such that \[ \large a=ck \]. Combining both equations we have \[ \large 1=ax_0+by_0=(ck)x_0+by_0=c(kx_0)+by_0 \] then g.c.d.(c,b)=1